How many pounds of candy selling for $5.80 per pound should be mixed with 7 punds of candy selling for $2.40 per pund to obtain a mixture for $3.42 per pound?
Answer by longjonsilver(2297)

this is an easy question to solve, once you get past all the english.

First let x = amount of candy required (at £5.80).

so, we have candy1 + candy2 = (mixture)...that is all there is to it!

so we have 5.80x + 7(2.40) = (7+x)*3.42
5.80x + 16.80 = 23.94 + 3.42x
2.38x = 7.14
x = 3

so, we need 3lb of the first candy.

jon.